An object of weight 120N vibrates with a period of 4.0s when hang from a spring. Calculate the force per unit length of the spring [g=10m/s, π=3.142]
1
Expert's answer
2019-05-03T08:56:01-0400

The period of oscillating spring is

"T=2\pi \sqrt{\frac{m}{k}}"


or

"T=2\pi \sqrt{\frac{P}{kg}}"

From this we have force per unit length

"k=\frac{4\pi^2P}{T^2g}"

Calculating

"k=\frac{4\cdot (3.142)^2\cdot 120}{10\cdot 4^2}=29.6 (\frac{N}{m})"